OpenStudy (anonymous):

Find the derivative of s(q)=3cosqsinq

OpenStudy (anonymous):

\[(fg)'=f'g+g'f\] with \[f(x)=\cos(x), f'(x)=-\sin(x), g(x)=\sin(x), g'(x)=\cos(x)\] make a direct substitution

OpenStudy (anonymous):

What do you do with the 3? Anything or does it become zero?